The sum of the ages of Petra and her mother is 48. Her mother is 9 years more than twice as old as Petra. How old are Petra and her mother?
Answer by ikleyn(52781)

.
Petra + Mother        = 48


Petra + (2*Petra + 9) = 48


P + (2P +9) = 48


3P = 48 - 9


3P = 39  ====>  P = 39/3 = 13.


Answer.  Petra is 13 years old;  the mother is 2*13+9 = 35 years old.

Solved.